### v2

# Learning Object State Changes in Videos: An Open-World Perspective

Abstract
Object State Changes (OSCs) are pivotal for video understanding. While humans can effortlessly generalize OSC understanding from familiar to unknown objects, current approaches are confined to a closed vocabulary. Addressing this gap, we introduce a novel open-world formulation for the video OSC problem. The goal is to temporally localize the three stages of an OSC—the object’s initial state, its transitioning state, and its end state—whether or not the object has been observed during training. Towards this end, we develop V ID OSC, a holistic learning approach that: (1) leverages text and vision-language models for supervisory signals to obviate manually labeling OSC training data, and (2) abstracts fine-grained shared state representations from objects to enhance generalization. Furthermore, we present HowToChange, the first open-world benchmark for video OSC localization, which offers an order of magnitude increase in the label space and annotation volume compared to the best existing benchmark. Experimental results demonstrate the efficacy of our approach, in both traditional closed-world and open-world scenarios. [^1] Project webpage: https://vision.cs.utexas.edu/projects/VidOSC/.
1 Introduction
Figure 1: Top: The video OSC objective is to temporally localize an object’s three states (i.e., initial, transitioning, end). Bottom: OSCs naturally exhibit a long tail. Certain OSCs, such as melting butter or marshmallow, are frequently showcased in instructional videos while others like melting jaggery might be rarely seen. We introduce an innovative open-world formulation that requires extrapolating to novel objects never encountered during training.
In video understanding, the study of objects primarily revolves around tasks like recognition [16], detection [24], and tracking [62, 8], with the assumption that objects maintain a consistent visual appearance throughout the video. Yet, objects in video are often dynamic. They can undergo transformations that change their appearance, shape, and even topology. For example, a pineapple goes from whole to peeled to sliced, or wood is carved into a new shape.
Object State Changes (OSCs) [15, 22, 2, 17, 50, 51, 52, 47, 64, 60] add a critical dimension to video understanding. On the one hand, they provide insights into human actions—observing a piece of metal being shaped into a hook, for instance, implies the action of bending; observing an egg shell go from whole to broken implies the action of cracking. On the other hand, OSCs are essential for assessing goal completion, in a way that is invariant to the specific procedure used. For instance, the readiness of cake batter signifies the completion of a mixing task, regardless of whether it was stirred by hand or a mechanical mixer. These core functionalities are instrumental for various real-world applications [20, 21, 13, 12, 6, 5, 66], ranging from AR/VR assistants [42] that guide users through complex tasks by monitoring object states, to robotic manipulation [55, 9] where understanding the state of objects is critical for task planning and failure recovery.
However, the development of video OSC is still at a primitive stage. Existing approaches [2, 50, 51, 15, 1, 46, 58] assume a closed vocabulary, where each state transformation is associated with a limited set of objects—often just 1 or 2. For example, the concept of “melting" might be limited to familiar items like butter and marshmallow. Consequently, the learned models are only capable of identifying state changes for objects observed during training and stumble when presented with novel objects. In contrast, in real-world situations a single state transition like “melting" can be linked with a plethora of objects—some ubiquitous and others rare. Importantly, there is an intrinsic connection that threads these OSC processes together. As humans, even if we have never encountered an unusual substance (e.g., jaggery) before, we can still understand that it is experiencing a “melting” process based on its visual transformation cues (see Fig. 1, bottom).
In light of these limitations, we propose a a novel open-world formulation of the video OSC problem. We formally characterize OSCs in videos in terms of three states that must be temporally localized: initial, transitioning, and end (see Fig. 1, top). In our open-world setting, there are known objects and novel objects. The model is only presented with known objects during training (e.g., frying chicken, frying onions). Then, it is evaluated on both known and novel objects (e.g., frying cauliflower). In addition to this fundamental open-world generalization, we also tackle a more comprehensive notion of the transitioning state. Specifically, our transitioning states encapsulate not only action-induced modifications to the object (e.g., peeling), but also passive transformations the object undergoes without human interference (e.g., melting) and edited-out transformations (e.g., the cake goes from raw to baked even if baked off camera). Though largely overlooked in existing approaches [15, 58, 1, 46], these cases are both common and technically interesting, since they force a model to reason about the effects of the core state change rather than search for evidence of a human actor carrying out the action.
Armed with this new problem formulation, we propose a holistic video OSC learning approach, anchored by two innovative ideas. First, we explore text and vision-language models (VLMs) for supervisory signals during training. We pioneer the use of textual state descriptions to generate a long tail of object state pseudo-labels, by leveraging the remarkable capabilities of VLMs [44, 23, 63] and large language models (LLMs) [53, 54, 40]. This strategy eliminates the need for exhaustive label collection for training data, facilitating large-scale model training. [^2] Note that at test time, our model requires only the video, with no need for additional text, ensuring utmost flexibility and applicability. Second, to confront the open-world challenge, we propose object-agnostic state prediction, consisting of three key techniques: a shared state vocabulary to unify the label space across objects sharing the same state transition (e.g., melting butter and melting marshmallow), temporal modeling to grasp the progression of state changes over time, and object-centric features that better represent the objects during an OSC. These designs equip our model with the ability to generalize state understanding from known objects to novel ones. We term our approach VidOSC.
Complementing this, we present a large-scale real-world dataset HowToChange. Sourced from the HowTo100M collection [36], it sets a new benchmark in the field of unprecedented scale and an authentic long-tail distribution, setting it apart from the constrained scope and closed-world setting of earlier benchmarks (see Table 1). Finally, experimental results demonstrate the efficacy of VidOSC, surpassing the state-of-the-art in both traditional closed-world and novel open-world scenarios by a great margin.
2 Related Work
Object State Changes Image-based methods explore the compositional nature of objects and their attributes, including zero-shot recognition of unseen combinations [37, 39, 22, 43, 33, 38, 41, 29], but do not consider the temporal progression of OSCs in video, which brings new challenges. Video-based methods develop human-centric models that leverage state change cues to facilitate action recognition  [15, 58, 1, 46], or explore joint discovery of object states and actions  [2, 50, 51]. Notably, all the existing methods assume a closed world, recognizing only the OSC categories (“known" objects) seen during training. Our approach distinguishes itself in three crucial ways: (1) we introduce a novel open-world333also called “unseen compositions” in object-attribute learning [37, 34]. formulation, where the objects encountered during evaluation can be entirely unseen during training; (2) we adopt an object-centric perspective, allowing for scenarios where OSCs occur with no observable human actions in the video; and (3) we propose to utilize the text modality and VLMs as supervisory signals, greatly scaling up model training and boosting performance.
Early datasets for video OSC capture a limited array of OSCs in fewer than 1,000 videos  [2, 31]. The more recent ChangeIt dataset [50] marks an advance in dataset scale (34K videos), yet is still restricted to a small OSC vocabulary of 44 total object-state transitions. It lacks the scope to adequately test unseen objects, since most state changes coincide with only 1 or 2 objects across all the videos (see Table 1). Other datasets explore different aspects of video OSCs, including object state recognition [47], point-of-no-return (PNR) frame localization [17], and object segmentation [52, 64], but they lack any temporal annotations of fine-grained OSC states needed for our task’s evaluation. Our HowToChange dataset increases the OSC vocabulary space by an order of magnitude and allows for the first time rigorous study of the open-world temporal OSC challenge.
Vision and Language LLMs [53, 54, 40] have revolutionized various research fields with their exceptional performance across diverse applications. Building on this momentum, the use of web-scale image-text data has emerged as a powerful paradigm in computer vision, with powerful VLMs now advancing an array of image tasks, including zero-shot classification [44], detection [18], segmentation [32] and visual question answering [27, 28]. Similarly, joint video-text representation learning [36, 30, 67, 3] has been advanced by multimodal datasets like HowTo100M [36] and Ego4D [17], which offer large-scale collections of videos paired with their corresponding narrations. These datasets have also facilitated tasks like step discovery [11], localization in procedural activities [35] and long egocentric videos [45]. The vision-language multimodal cycle consistency loss proposed in [14] helps discover long-term temporal dynamics in video. In line with these developments, we propose to leverage the text accompanying instructional videos as well as existing high-performing VLMs to provide supervisory signals that guide the training of our video OSC model and allow learning for the open world.
Learning in an Open World The open world setting has received increasing attention, predominantly within the image domain for object recognition [48, 4], detection [10, 25, 18, 65] and object-attribute compositionality [37, 39, 43, 33, 38, 41, 29]. In the video domain, compositionality is advanced by the Something-Else dataset [34], where the training combinations of verbs and nouns do not overlap with the test set, sparking work on the dynamics of subject-object interactions [34, 47, 7]. More recent efforts leverage VLMs for zero-shot action recognition  [57, 26, 7]. Concurrent work explores video object segmentation with state-changing objects [64] and recognition of novel object-state compositions for food chopped in different styles [47]. Despite these advances, temporal video OSC understanding in the open world remains unexplored. Our open-world formulation requires generalizing the temporal localization of fine-grained object states from known to novel objects.
3 Approach
Figure 2: Our proposed VidOSC framework: (a) Mining for OSC examples (Sec. 3.3): We leverage ASR transcriptions paired with videos and the capabilities of LLM to automatically mine OSC examples. (b) Pseudo Label Generation (Sec. 3.3): We utilize textual state descriptions and a VLM for supervisory signals during training; (c) Model Training (Sec. 3.2): We develop a video model for object-agnostic state prediction. (d) Model Testing (Sec. 3.1): We propose an open-world formulation, evaluating on both known and novel OSCs. Notably, while we employ the text modality to guide model training, our model is purely video-based and requires no text input at the test phase, ensuring maximum flexibility and applicability. Ground truth for the test set is manually annotated.
We present the VidOSC framework for learning video OSCs, detailing the open-world formulation (Sec. 3.1), model design for object-agnostic state prediction (Sec. 3.2), and text-guided training scheme (Sec. 3.3).
3.1 Video OSC in the Open World
We begin by formally defining an object state change (OSC) in videos as a visually detectable transformation, where an object experiences a change that is not easily reversible, in line with [17]. We characterize an OSC category as an object combined with a specific state transition, such as “chicken + shredding”, and delineate an OSC process through three distinct states: initial (precondition state), transitioning, and end (postcondition state). [^4] Due to real-world data variation, some videos may lack one of the OSC states, and there may be multiple segments in a video corresponding to the same state. Our formulation accounts for all these scenarios. It is essential to highlight that we take an object-centric perspective: the “transitioning state” accounts for instances where the video depicts an active action applied to the object, as well as scenarios where the object undergoes a passive transformation absent of human intervention (e.g., melting, drying).
Given a video in which an object may be changing state, the objective is to temporally localize each of the three OSC states. Consistent with prior work [50, 51], we formulate the task as a frame-wise classification problem. Formally, a video sequence is represented by a temporal sequence of $T$ feature vectors $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}\}$, where $T$ denotes the video duration. The goal is to predict the OSC state label for each timepoint, denoted by $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\cdots,\mathbf{y}_{T}\}$, $\mathbf{y}_{t}\in\{1,\cdots,K+1\}$, where $K$ is the total number of OSC states, and there is one additional category representing the background class not associated with any OSC state. [^5] Not all training videos may feature an OSC due to data collection noise, yet all evaluation videos are manually verified to include one.
Next, we propose an open-world problem formulation, capturing the intrinsic long-tail distribution observed in real-world scenarios. Consider $N$ state transitions, each paired with a set of associated objects, denoted by $\mathbf{O}^{n}=\{o_{1}^{n},o_{2}^{n},\cdots,o_{m(n)}^{n}\}$, for $n=\{1,2,\cdots,N\}$, where $m(n)$ denotes the number of objects linked with the $n$-th state transition. Within a specific state transition (like frying), certain objects in the set (such as chicken or onions) are frequently observed, while the combination of the same state transition with other objects (such as cauliflower) appear less often. Motivated by this inherent long-tail, we propose to split $\mathbf{O}^{n}$ into two distinct subsets: $\mathbf{O}_{\text{known}}^{n}$ covering the common combinations observed during training, and $\mathbf{O}_{\text{novel}}^{n}$ comprising the infrequent combinations, which are unseen during training due to their rarity. During inference, the model is evaluated on the entire object set $\mathbf{O}^{n}$ for a comprehensive reflection of the open-world setting.
3.2 Object-Agnostic State Prediction
To address the open-world challenges and ensure robust generalization to novel objects, VidOSC integrates three key techniques: (1) a shared state vocabulary that consolidates the state label space; (2) temporal modeling within the video OSC model design; and (3) object-centric features.
Shared State Vocabulary The inherent link among different OSC categories is overlooked in prior work. Common practice involves developing separate models for each OSC category [2, 50] (e.g., one model exclusively for melting butter and another for melting chocolate), or using one single model that still treats every OSC as a distinct entity in the label space [51] (e.g., considering melted butter and melted chocolate as two separate end states). Such a formulation results in an extensive label set with $K=3\times\sum_{n=1}^{N}m(n)$, where 3 denotes the number of OSC states (i.e., initial, transitioning and end) and $\sum_{n=1}^{N}m(n)$ is the total number of OSC categories. This compartmentalized view can inadvertently hinder the model’s generalization ability. Yet, on closer inspection, states across varied objects are intrinsically related. For instance, the “melting” principle remains consistent, even when applied to visually distinct objects like butter and chocolate. This motivates us to group OSC labels that share the same state transition as one, and train a single model per state transition. [^6] See Supp. for the multi-task model variant, where we develop one unified model for all state transitions. By adopting object-agnostic OSC labels, we encourage the model to learn state representations shared by visually different objects, and thus facilitate the transfer from known to novel objects.
Temporal Modeling Recognizing that state changes unfold over time, it is important to capture the temporal dynamics in videos. An object’s state at any frame, be it initial, transitioning, or end, is often best understood in the context of preceding and succeeding frames. Contrary to prior works [2, 50, 51] that rely on isolated frame-wise modeling without considering the temporal progression of video OSCs, we address this gap by proposing a temporally-aware model design. As illustrated in Fig. 2(c), we adopt an encoder-decoder architecture. For the encoding phase, input video features $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}\}$ are first projected using an MLP $f_{\text{project}}$, and augmented with sinusoidal positional embeddings, $\mathbf{Z}_{\text{pos}}$. Subsequently, a transformer encoder [56] $f_{\text{transformer}}$ is adopted to capture the temporal dynamics among these features, yielding $\mathbf{Z}=f_{\text{transformer}}(f_{\text{project}}(\mathbf{X})+\mathbf{Z}_{\text{pos}})$. For the decoding phase, a MLP decoder $g$, maps these temporally-aware hidden representations to OSC state predictions: $\tilde{\mathbf{Y}}=g(\mathbf{Z})$. See Sec. 5 for architecture and training details. This design ensures that when predicting the state for a frame, the model has assimilated temporal context from the entire sequence. Essentially, our model exploits the fact that the dynamics of the state change has greater invariance to the object category than how the object looks, emphasizing temporal object transformations over objects’ static appearances, and thereby enhancing generalization to novel OSCs.
Object-Centric Features Finally, we discuss how to better represent “object” features in the problem. In many scenarios, due to camera placement and framing, the object going through a state transition might only occupy a small portion of the frame, surrounded by other visual elements such as background, people, or bystander objects. Recognizing this challenge, we introduce an enhancement to our model’s input features $\mathbf{X}$ to emphasize the object of interest. To be specific, we leverage an off-the-shelf detector [49] to identify the active object (i.e., the object being manipulated) region at each timepoint $t$, yielding feature $\mathbf{x}_{t}^{obj}$ that centers on the object (see bounding boxes in Fig. 2 (c)). The input feature is then constructed as a concatenation of the original global feature and the localized object-centric feature, i.e., $\mathbf{X}=\{[\mathbf{x}_{t},\mathbf{x}_{t}^{obj}]\}_{t=1}^{T}$. By emphasizing the object in this manner, our model is better positioned to discern the intricate state changes and provide more informed predictions.
3.3 Text and VLMs as Supervision
To scale up training and ensure broad generalization, we propose a novel training pipeline that leverages LLMs and VLMs for OSC mining and pseudo-labeling.
Mining for OSC examples Utilizing a vast collection of “how-to" instructional videos as the training source, we develop an automated mining process to capture the rich, real-world variability of OSCs. The motivation is that instructional videos usually have accompanying Automatic Speech Recognition (ASR) transcriptions that offer valuable OSC cues. For example, a speaker mentioning, “so just get your rotisserie chicken and shred it up” suggests that the chicken may undergo a state transition of shredding in the video. Leveraging this fact, we employ LLAMA2 [54] to analyze ASR transcriptions for identifying candidate videos and their associated OSC categories. This text mining stage allows us to corral a long tail of OSCs, discovering likely state change terms—even if rare—from the free-form natural language narrations given by how-to instructors in the training videos. See Fig. 2 (a).
Pseudo Label Generation Next, to negate the need of manually labeling large-scale training data, we propose a novel pseudo-labeling pipeline facilitated by VLMs. From the identified OSC category (e.g., shredding chicken), we form three textual state descriptions for its initial, transitioning, and end states (e.g., whole chicken, shredding chicken, and shredded chicken). We then adopt both the vision and language encoder from a well-trained VLM (we experiment with CLIP [44] and VideoClip [61]) to compute the cross-modal similarity between every frame in training video and the three state descriptions, producing a score matrix $\mathbf{S}\in\mathbb{R}^{T\times 3}$. The pseudo label $\hat{y}_{t}$ at timepoint $t$ is then assigned based on this score matrix:
$\hat{y}_{t}=\begin{cases}\text{Background}&\text{if }{\sum}(\mathbf{S}[t,:])<\tau\\ \text{Initial}&\text{elif }\mathbf{S}[t,0]-\mathbf{S}[t,1]>\delta\text{ and }\mathbf{S}[t,0]-\mathbf{S}[t,2]>\delta\\ \text{Transitioning}&\text{elif }\mathbf{S}[t,1]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,1]-\mathbf{S}[t,2]>\delta\\ \text{End}&\text{elif }\mathbf{S}[t,2]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,2]-\mathbf{S}[t,1]>\delta\\ \text{Ambiguous}&\text{otherwise}\end{cases}$
where $\delta$ is the threshold that separates states and $\tau$ is the threshold that differentiates between states and background. Essentially if the VLM scores some state more strongly than the other two—and the cumulative confidence score of all states is high—then it is adopted as the pseudo label. Otherwise it is omitted as ambiguous. See Fig. 2(b).
To further refine these labels, we enforce a causal ordering constraint. Given the inherent progression of OSCs, the anticipated order is initial states followed by transitioning states, and finally, the end states. Any frame whose pseudo label does not respect this natural progression is re-assigned to the ambiguous category. Our training employs a cross-entropy loss between pseudo label $\hat{y}_{t}$ and the corresponding model prediction $\tilde{y}_{t}$, with ambiguous frames excluded to maintain clarity and distinction among the state labels. See Section 4.2 in Supp. for detailed pseudo label analysis.
4 The HowToChange Dataset
Existing OSC datasets fall short in capturing the open-world’s diversity and long-tail distribution of OSCs. To address this gap, we introduce HowToChange. It encompasses varied state transitions coupled with a diverse range of objects, providing an authentic reflection of their real-world frequency—from the commonplace to the rare.
Figure 3: Ground truth annotation distribution across 20 state transitions (top) and 134 objects (bottom) in HowToChange (Evaluation). In line with our open-world formulation, annotations cover a diverse range of object-state transition combinations, categorized into known and novel OSCs.
Data Collection The HowTo100M collection [36] contains a wealth of instructional videos that often feature OSCs and is particularly suitable for this task. We specifically focus on the HowTo100M Food & Entertaining category because (1) it constitutes a third of the entire HowTo100M videos, (2) cooking tasks offer a wealth of objects, tools, and state changes, providing an excellent test bed for open-world OSC, and (3) in cooking activities a single state transition can often be associated with a varied range of objects, opening the door to learning compositionality. (Note, we also experiment with non-cooking domains below.) We process a total of 498,475 videos and 11,390,287 ASR transcriptions with LLAMA2. From the responses, we identify the most frequently seen state transitions and objects associated with them to establish an OSC vocabulary, resulting in 134 objects, 20 state transitions, and 409 unique OSCs. The number of objects associated with each state transition ($m(n)$) spans from 6 for “zesting” to 55 for “chopping”. The state transitions applied to each object vary from 1 to 15, with onions being the most versatile. In total, we identify 36,075 videos for the training set of HowToChange, with an average duration of 41.2 seconds.
Data Splits In line with our open-world formulation, we divide the 409 identified OSCs into two disjoint subsets based on their frequency of occurrence. Within each state transition, we categorize the top 75% frequent objects as known and the bottom 25% as novel. This yields a total of 318 known OSCs that are seen during training and testing, spanning 20 state transitions associated with 122 objects, and 91 novel OSCs that are only seen during testing, encompassing the same transitions across 58 objects.
Ground Truth Label Collection (Evaluation Set) To facilitate thorough evaluation, we obtain manual annotations for a subset of 5,424 videos from the collected dataset. [^7] Our training is purely guided by a VLM and requires no ground truth labels. The annotations are reserved exclusively for evaluation. The annotation workflow is as follows: each annotator is presented with a video segment along with an OSC category that was previously identified by LLAMA2. The annotator has the option to reject the video segment if it does not contain the specified OSC. Otherwise, they label the time ranges corresponding to the initial, transitioning, and end states of the OSC. Adhering to our object-centric emphasis, annotators are instructed to label based on the visual changes of the object, rather than human-centric actions, and exclude time ranges where the object of interest is not visible, ensuring clean and focused temporal labels. Fig. 3 provides the distribution of annotated videos. On average, we collect 271 annotations per state transition, with a video duration of 41.5 seconds, and 12.9% of videos belong to the novel category. The entire annotation process required around 1,507 hours by 30 professional annotators.
 Datasets # Obj # ST # OSC ObjPer # Videos GT Label?
Alayrac et al. [2] 5 6 7 1.2 630
Task-Fluent [31] 25 14 32 2.3 809
ChangeIt (Training) [50] 42 27 44 1.6 34,428
ChangeIt (Evaluation) [50] 42 27 44 1.6 667
HowToChange (Training) 122 20 318 15.9 36,075
HowToChange (Evaluation) 134 20 409 20.5 5,424
Table 1: Comparison with existing video datasets focusing on object states. ‘Obj’ and ‘ST’ represent objects and state transitions, respectively. ‘ObjPer’ denotes the average number of objects associated with each state transition; higher values indicate more need to generalization across objects. We present the first open-world benchmark for temporal video OSC, with an order of magnitude increase in OSC vocabulary and annotation volume.
Dataset Comparison Table 1 compares HowToChange with existing video datasets on temporal OSC understanding. HowToChange offers an unprecedented scale—with 9.3x more OSC categories and 8.1x more annotated video instances compared to the previous largest collection [50]. Furthermore, notably, in prior datasets, each state transition is typically coupled with 1 or 2 objects, preventing subsequent models from generalizing to new objects, as we will see in results. In contrast, HowToChange pioneers the open-world formulation, presenting a much broader range of objects associated with each state transition—from 6 to 55 objects per state transition, and averaging 20. This facilitates the development of models with generalized OSC understanding. Please see Supp. for full data collection and annotation details.
5 Experiments
 Method ChangeIt ChangeIt (open-world) HowToChange
State Prec.@1 Action Prec.@1 State Prec.@1 Action Prec.@1 F1 (%) Prec (%) Prec.@1 (%)
known novel known novel known novel known novel known novel
CLIP [44] 0.30 0.63 0.29 0.29 0.71 0.70 26.9 25.4 27.3 26.6 47.5 47.5
VideoCLIP [61] 0.33 0.59 0.25 0.24 0.62 0.55 36.6 34.3 39.7 38.5 48.3 44.8
InternVideo [59] 0.27 0.57 0.29 0.25 0.60 0.61 29.9 29.5 31.4 30.8 46.9 46.3
LookForTheChange [50] 0.35 0.68 0.36 0.25 0.77 0.68 30.3 28.7 32.5 30.0 37.2 36.1
MultiTaskChange [51] 0.49 0.80 0.41 0.22 0.72 0.62 33.9 29.9 38.5 34.1 43.1 38.8
VidOSC (ours) 0.57 0.84 0.56 0.48 0.89 0.82 46.4 43.1 46.6 43.7 60.7 58.2
Table 2: Results on ChangeIt, ChangeIt (open-world), and HowToChange. VidOSC outperforms all approaches in both closed-world and open-world scenarios, across known and novel OSCs.
Figure 4: Top-1 frame predictions given by VidOSC for the initial, transitioning, and end states, on ChangeIt (open-world) (first 2 rows) and HowToChange (last 2 rows). VidOSC not only accurately localizes the three fine-grained states for known OSCs, but also generalizes this understanding to novel objects, such as cauliflower and capsicum, which are not observed during training.
Datasets In addition to our new HowToChange dataset, we also evaluate on ChangeIt [50] due to its expansive data scale (34K videos spanning many activities) and high relevance to our task. Beyond the conventional split of ChangeIt, we propose a new split tailored to our open-world formulation. Specifically, from the 44 available OSC categories in ChangeIt, we concentrate on the 25 categories where each state transition is paired with more than one object. With those, we form the “ChangeIt (open-world)" subset that comprises 8 state transitions and 25 objects. Within each state transition, objects are randomly divided into known and novel categories, yielding 13 known OSC categories for training and 12 novel OSC categories exclusively reserved for evaluation. A detailed breakdown of this split can be found in Supp. To sum up, our evaluation encompasses: ChangeIt; ChangeIt (open-world); and HowToChange, offering a comprehensive setup in both closed-world and open-world scenarios.
Evaluation For ChangeIt and ChangeIt (open-world), we adhere to the original dataset’s evaluation protocol [50], reporting action and state precision@1 as the evaluation metrics. For our dataset, besides precision@1, which evaluates a single frame for each state within a video, we advocate for the use of F1 score and precision over all frames to ensure a more holistic evaluation.
Baselines We compare our approach with four baselines across two categories: (a) self-supervised approaches on identifying object states enforced by the causal ordering constraint: LookForTheChange [50] trains a dedicated model for each OSC category, while MultiTaskChange [51] evolves this into a multi-task approach [^8] To ensure a thorough evaluation, we train both single-task and multi-task variants of our approach. See Supp. for a detailed discussion. , catering to several OSCs concurrently; (b) zero-shot VLMs: image-based CLIP [44], video-based VideoCLIP [61] and InternVideo [59]. All baselines in (a) use the same training data as our model, whereas the zero-shot models (b) are directly evaluated on the test set with no training.
Implementation Videos are sampled at one frame per second. Each one-second video segment gets assigned to an OSC state label, and is encoded by InternVideo [59], a general video foundation model (which is kept frozen for training efficiency). Our video model consists of a 3-layer transformer with a hidden dimension of 512 and 4 attention heads as the encoder and a 1-layer MLP as the decoder. Consistent with prior work [50, 51], our model predicts the OSC state label (i.e., initial, transitioning, end state, or background) for a video, assuming the video OSC category is known (say from a recognition model’s output or a user’s specific query). While the standard output does not include the state transition name, our multi-task version detailed in Supp. is capable of this.
Main Results Table 2 presents results on all three datasets. VidOSC outperforms all approaches in both traditional closed-world and our newly introduced open-world settings. The large performance gains—as much as 9.6% jumps in precision vs. the next best method—underscore the effectiveness of two pivotal components in VidOSC. First, its use of text and VLMs for supervisory signals during training: particularly on novel OSCs , VidOSC extracts meaningful cues from the video modality to refine pseudo labels from VLMs, and ultimately surpasses the VLM baselines. Second, our model for object-agnostic state prediction, designed for the open-world context, effectively narrows the gap between known and novel OSCs. For instance, on ChangeIt (open-world), MultiTaskChange [51] experiences a 19% decline in state precision@1 while VidOSC has only a 8% drop. Note that we observe a more pronounced performance drop of all approaches on ChangeIt (open-world) than on HowToChange due to its limited number of objects available per state transition. These results also point to the potential for future development in bridging the known and novel OSC performance gap.
Qualitative Results Fig. 4 presents VidOSC’s top-1 frame predictions on ChangeIt (open-world) and HowToChange, across both known and novel OSCs. Notably, despite never seeing objects such as cauliflower and capsicum during training, VidOSC effectively leverages visual state change cues and correctly localizes the three states of these objects going through OSCs, demonstrating its strong generalization capability. For a more holistic view, Fig. 5 compares VidOSC’s frame-by-frame predictions with all baselines for a given test video. VidOSC provides temporally coherent predictions, smoothly progressing through the OSC states in the natural order (i.e., from initial to transitioning then end). In contrast, baseline approaches often yield fragmented and inconsistent predictions, indicating a lack of understanding of the video’s global temporal context, primarily due to their reliance on frame-wise modeling. See Supp. and Supp. video for more qualitative examples and VidOSC’s interpretability on object relations.
Figure 5: Comparison of model predictions across a test video depicting the OSC of “slicing shallot” on HowToChange. The x-axis represents temporal progression through the video. VidOSC gives temporally smooth and coherent predictions that best align with the ground truth, significantly outperforming baselines in capturing the video’s global temporal context.
 Shared Temporal Object Prec.@1 (%)
State Modeling Centric known novel $\Delta$
58.5 53.3 5.2
52.9 48.2 4.7
59.8 56.7 3.1
60.7 58.2 2.5
Table 3: Ablation Study. $\Delta$ denotes the performance gap between known OSCs and novel OSCs.
Ablation To further dissect the performance gains brought by our three model design techniques (i.e., shared state vocabulary, temporal modeling, and object-centric features), we conduct an ablation study, removing one component at a time. Table 3 confirms the essential role of each element. A shared state vocabulary is particularly crucial in the open-world context, as its absence increases the gap between known and novel OSCs from 2.5% to 5.2%. Furthermore, temporal modeling provides a substantial performance boost, and object-centric features offer further gains. See Supp. for an additional analysis of VidOSC’s performance with different pseudo labels.
Frame Retrieval VidOSC learns features that accurately characterize the evolution of an OSC process. To illustrate, we consider a novel frame retrieval setting. Within the HowToChange test set, when presented with a query video featuring a novel OSC at its initial state, VidOSC seeks the most similar and most contrasting video from a pool of candidates at their end states. The frame triplets with the smallest and largest feature distance are shown in Fig. 6. Remarkably, the retrieved nearest post-OSC frames correspond to the query’s anticipated state transition trajectory, despite the object and state gap. Conversely, the furthest frames exhibit end states of markedly different objects. These results further lend support to VidOSC’s capability to understand the evolution of an OSC process, even for novel objects it has never encountered during training.
Figure 6: Frame retrieval on the HowToChange test set. Given a query frame showcasing a novel OSC at its initial state, VidOSC retrieves the nearest and furthest frame among all known OSCs at their end states. The closest post-OSC frame follows the transition trajectory of the query while the furthest post-OSC frame depicts a substantially different object, demonstrating VidOSC’s capability to generalize the OSC progression for novel objects.
6 Conclusion
This work aims at a comprehensive exploration of video OSCs, with a novel open-world formulation. To address the challenges, we leverage text and VLMs to assist the training of a video OSC model at scale and design three modeling techniques to achieve object-agnostic state prediction for better generalization to novel OSCs. Furthermore, we present the most expansive video OSC dataset collection HowToChange, which echoes the natural long-tail of state transitions coupled with varied objects, fostering a realistic representation of real-world scenarios. As for future work, we will consider extending VidOSC to video sequences featuring concurrent OSC processes, and integrating spatial understanding of OSC within our open-world framework.
Acknowledgements: UT Austin is supported in part by the IFML NSF AI Institute. KG is paid as a research scientist by Meta.